OpenStudy (anonymous):

solve this system of equation using subtraction x + 3y = 2 2x - 3y = 22

OpenStudy (goformit100):

take all the terms on one side and subtract, the equations

OpenStudy (zzr0ck3r):

the first thing I did was -2(x+3y=2)

OpenStudy (ganpat):

multiply equation 1 by 2.. So, 2x + 6y = 4 ----- eqtn 1 2x - 3y = 22 ---- equation 2 subracting... 9y = -18 so y = -3.. and x = -7..

OpenStudy (goformit100):

i was correct

OpenStudy (ganpat):

sorry, x =11..